# Code-switching Speech Recognition Under the Lens: Model- and Data-Centric Perspectives

### v1

@@ IV-B Experiment setup and model configuration
Experiments with models trained from scratch, such as Conformer, were conducted using ESPNet [35]. For training on the SEAME and ASRU datasets, we adopted the hyperparameter settings from [19] for Conformer and from [53] for Branchformer and ConExtBiMamba. Pre-training on monolingual datasets followed the hyperparameters specified in the LibriSpeech recipe provided by ESPNet. The resulting model is referred to as Conformer-L, reflecting its larger hidden dimension of 512, in contrast to the 256-dimensional hidden size used in the Conformer models trained solely on SEAME or ASRU data. This setup is also employed for Conformer-L Mix, which is pre-trained on a mixture of code-switching and monolingual datasets. Inference was performed using the average of the weights from the ten best-performing models on the development sets, with a beam size of five.
Experiments with Whisper models were conducted using Whisper-small [9]. Fine-tuning employed the AdamW optimizer, with the learning rate linearly warmed up from 0 to $1\times 10^{-5}$ over 10,000 update steps, followed by a linear decay schedule. Training used a batch size of eight with gradient accumulation over two steps. For inference, we adopted greedy decoding with model weights averaged from the three best-performing checkpoints on the development set. Language prompts <<zh>> and <<en>> were applied to the ASRU and SEAME datasets, respectively. LoRA with a rank of 24 was integrated into the feed-forward networks of Whisper [54]. Following the Bi-encoder mechanism [12], which is described in equations (4)–(6), Bi-LoRA was pre-trained separately on LibriSpeech and ASRU Mandarin, and subsequently fine-tuned on either ASRU or SEAME, with each LoRA configured to a rank of 12.
Speech synthesis was performed using the open-source TTS model CosyVoice2 [7], which was fine-tuned on the target domain data before synthesis. Speech samples from different datasets as prompt speech were used to simulate various accents: LibriSpeech for American English, ASRU for Mandarin spoken in mainland China, and SEAME for Southeast Asian-accented English and Mandarin.
@@ IV-C Performance measurement
In addition to the ASR performance, we adopt an LLM-as-a-judge method to evaluate the quality of texts generated by LLMs. Specifically, GPT-4o assigns a score from 1 to 10 to each sample with reference to the criteria described in Table VIII. The final score is computed as the average across all text samples. We also evaluate the similarity in language confusion between real text and LLM-generated text by comparing their CMI.
@@ V-A Comparison of model-centric methods
Params. test $\downarrow$ $\text{dev}_{\texttt{man}}$ $\downarrow$ $\text{dev}_{\texttt{sge}}$ $\downarrow$
Transformer [57] 29.8 M 13.1 17.7 24.5
@@ V-B Impact of Pre-training
We also explored the impact of pre-training strategies on CS-ASR performance and presented the results in Table IV. Training a Conformer-L model directly on the mix of monolingual data and code-switching data obtains higher performance than sequentially pre-training on monolingual data and fine-tuning on code-switching data. This suggests that CS-ASR systems benefit from both cross-lingual and monolingual knowledge during training. While incorporating code-switching data in the training data moderately degrades monolingual ASR performance, it results in higher overall performance than both the pre-training and the fine-tuning strategies. The pre-trained Whisper-small model demonstrates an inherent ability to handle code-switching due to its large-scale multilingual pre-training. Fine-tuning the Whisper-small model on the ASRU leads to higher performance on both in-domain ASRU Mandarin and code-switching test data, while significantly degrading the performance on the LibriSpeech test-clean set.
@@
TABLE V: Comparison of fine-tuning the Whisper-small model on real, synthesized speech, and data augmented with synthesized speech. Here, the target texts and prompt samples are sourced from the original ASRU dataset. Real and synthesized data are denoted as “Real” and “Syn.”, respectively
@@ V-C TTS with original text and speaker variety
Results indicate that augmenting training with synthesized speech and original text improves performance on SEAME but degrades performance on ASRU. When augmented with synthesized data three times the size of the original training set, the fine-tuned model achieves the best WER of 12.7% and 17.7% on the two SEAME test sets, respectively. However, this data augmentation strategy leads to performance degradation for ASRU. This discrepancy can be attributed to differences in textual complexity between the SEAME and ASRU datasets.
@@
V-D Impact of accents and language confusion on ASR
Since the $\text{dev}_{\texttt{man}}$ set has a CMI closer to the SEAME training data, the performance gap between $\text{dev}_{\texttt{man}}$ and $\text{dev}_{\texttt{sge}}$ presented in Table III suggests that the model performs worse on data with a higher degree of code mixing. To further examine how accent and language confusion in code-switching data affect ASR performance, we synthesized speech using reference prompts and target texts from multiple datasets, respectively. The resulting synthetic data were then used to fine-tune the Whisper-small model, and performance was compared across conditions.
To simulate accent mismatch, we paired text samples from the target dataset with prompt speech samples from a different dataset while keeping the TTS model consistent with the target dataset. For example, using a TTS model fine-tuned on SEAME data to synthesize speech for SEAME text with LibriSpeech or ASRU prompts results in an accent mismatch. Conversely, the textual mismatch, primarily associated with syntactic and lexical characteristics, was simulated by pairing out-of-domain text with an in-domain prompt and TTS model.
@@ V-E Evaluation of the proposed SECT prompt
TABLE VII: Comparison of code-switching text synthesis with different LLMs and prompts, performance is evaluated with code-switching samples synthesized via 1500 ASRU Mandarin text samples, where the fine-tuned model is tested on the ASRU test set. The CMI of the ASRU training set is 17.0
Method ASR GPT-4o CMI
@@
DeepSeek-R1 w/ baseline prompt 22.9 6.8 32.2
DeepSeek-R1 w/ EZSwitch prompt [33] 22.2 6.0 24.6
DeepSeek-R1 w/ SECT prompt (ours) 19.2 7.0 17.3
w/o in-context learning examples 19.6 7.0 15.2
@@
We evaluated the effectiveness of the proposed SECT-guided text generation method on 1,500 synthesized text samples in terms of WER for ASR performance, GPT-4o scores, and CMI. ASR performance was evaluated by fine-tuning a Whisper-small model on speech synthesized from the SECT-generated texts before test. GPT-4o scores were obtained via an LLM-as-a-judge protocol, used together with CMI to assess the naturalness, grammatical correctness, and semantic preservation of the generated text samples.
A comparison between SECT-guided prompts and other prompting strategies is presented in Table VII, and the prompting strategies and LLM-as-a-judge evaluation are introduced in Table VIII. The results demonstrate that the SECT-guided prompting strategy achieves the best performance in terms of both WER and GPT-4o scores, and yields a CMI closer to that of the real ASRU code-switching text.
Moreover, we conducted ablation studies by comparing three prompting strategies: the baseline prompt, the SECT prompt without in-context learning, and the SECT prompt with in-context learning. Results confirm that the effectiveness of the SECT-based prompting strategies since it consistently outperforms the baseline regardless of whether in-context examples are included. Although incorporating in-context learning examples does not improve the GPT-4o scores, it leads to lower WER and a CMI value that more closely aligns with the real ASRU data. This implies that while ECT provides beneficial structural constraints for generating linguistically valid code-switching text, it may also impose rigid switching patterns. Incorporating in-context examples helps mitigate such rigidity, leading to more natural code-switching behavior.
@@ V-F TTS with text generated via SECT-prompted LLM
TABLE IX: Comparison between fine-tuning the Whisper-small model on ASRU data with different data augmentation methods. Text synthesis is performed via LLM guided by the proposed SECT prompts, and TTS is performed using target texts sourced from “Text source” and reference prompts sourced from ASRU training set. Real and synthesized data are denoted as “Real” and “Syn.”, respectively
@@
+ $ft$ ASRU + ASRU Mandarin - 193 + 100 8.9
+ $ft$ ASRU + ASRU Mandarin - 193 + 500 9.2
@@
+ $ft$ ASRU + Syn. w/ Real text 193 + 180 9.3
+ $ft$ Syn. text w/ Qwen-chat 352 18.4
+ $ft$ Syn. text w/ DeepSeek-R1 ASRU 509 13.8
+ $ft$ ASRU + Syn. text w/ DeepSeek-R1 Mandarin 193 + 100 8.5
+ $ft$ ASRU + Syn. text w/ DeepSeek-R1 193 + 247 8.3
+ $ft$ ASRU + Syn. text w/ DeepSeek-R1 193 + 509 8.4
@@
We also investigated the impact of varying the amount of synthesized speech–text pairs on CS-ASR performance. Results show that increasing the amount of pure Mandarin data during Whisper fine-tuning leads to performance degradation. This finding is consistent with the results presented in Table V, where data augmentation introduces limited code-switching patterns and results in lower performance. In contrast, incorporating synthesized code-switching speech–text pairs into fine-tuning yields moderate improvements. Notably, the highest ASR performance is observed when the fine-tuning data includes a balanced amount of real and synthetic code-switching data, suggesting that carefully curated synthetic data can effectively complement limited real-world code-switching resources. This also implies that the synthesized text data still falls short of real-world text in terms of naturalness and linguistic richness.
@@
VI Conclusion
In this paper, we investigated CS-ASR from both model- and data-centric perspectives, focusing on language confusion, accent bias, and data scarcity. On the model side, we compared language-specific processing and multi-task learning, providing insights into their trade-offs across training and inference. On the data side, we explored TTS-based augmentation and showed that syntactic and lexical characteristics have a greater impact on CS-ASR performance than accent. To further mitigate data scarcity, we proposed SECT, a prompting strategy that guides LLMs to generate linguistically valid code-switching text. Combined with TTS, SECT produces diverse training data and outperforms existing prompting strategies. Comparing model- and data-centric approaches, we found that algorithmic methods are more effective in low-confusion scenarios with clear matrix–embedded language structures, while TTS-based augmentation is beneficial across different levels of confusion. Simple TTS augmentation with speaker variation improves performance in highly confused settings, whereas creating new code-switching text is essential for low-confusion data where textual diversity is limited.

### v2

@@ IV-B Experiment setup and model configuration
Experiments with models trained from scratch, such as Conformer, were conducted using ESPNet [55]. For training on the SEAME and ASRU datasets, we adopted the hyperparameter settings from [27] for Conformer and from [67] for Branchformer and ConExtBiMamba. Experiments involving pre-training on monolingual datasets followed the hyperparameters specified in the LibriSpeech recipe provided by ESPNet. The resulting model is referred to as Conformer-L Mono. Conformer-L has a larger hidden dimension of 512, in contrast to the 256-dimensional hidden size used in the Conformer models. This setup is also employed for Conformer-L Mix, which was developed on a mixture of code-switching and monolingual datasets. Inference was performed using the average of the weights from the ten best-performing models on the development sets, with a beam size of five.
Experiments with Whisper models were conducted using Whisper-small [43]. Fine-tuning employed the AdamW optimizer, with the learning rate linearly warmed up from 0 to $1\times 10^{-5}$ over 10,000 update steps, followed by a linear decay schedule. Training used a batch size of eight with gradient accumulation over two steps. For inference, we adopted greedy decoding with model weights averaged from the three best-performing checkpoints on the development set. Language prompts <<zh>> and <<en>> were applied to the ASRU and SEAME datasets, respectively. LoRA with a rank of 24 was integrated into the feed-forward networks of Whisper [16]. Following the Bi-encoder mechanism [29], which is described in equations (4)–(6), Bi-LoRA was pre-trained separately on LibriSpeech and ASRU Mandarin, and subsequently fine-tuned on either ASRU or SEAME, with each LoRA configured to a rank of 12. This ensures a fair comparison between general LoRA and bi-LoRA adaptation methods while achieving computational efficiency.
Speech synthesis was performed using the open-source TTS model CosyVoice 2 [11], which was fine-tuned on the target domain data before synthesis. CosyVoice 2 is pre-trained on massive English and Mandarin data, providing rich prior bilingual knowledge. This is critical for English-Mandarin code-switching, as the model has already learned the joint distribution and phonetic transitions between these two languages. This prior knowledge ensures that the synthesized code-switching data is phonetically accurate, providing higher effectiveness for developing CS-ASR models. Speech samples from different datasets as prompt speech were used to simulate various accents: LibriSpeech for American English, ASRU for Mandarin spoken in mainland China, and SEAME for Southeast Asian-accented English and Mandarin.
@@ IV-C Performance measurement
In addition to the ASR performance, we adopted an LLM-as-a-judge method to evaluate the quality of texts generated by LLMs. Specifically, GPT-4o assigns a score from 1 to 10 to each sample with reference to the criteria described in Table VIII. The final score is computed as the average across all text samples. Human evaluation was conducted involving ten bilingual graduate students proficient in both English and Mandarin. The evaluators were instructed to assess the generated text samples for each method using the same scoring criteria as the LLM-as-a-judge framework. We also evaluate the similarity in language confusion between real text and LLM-generated text by comparing their CMI and perplexity (PPL). The PPL is calculated with a Transformer-based language model (LM) developed on the training set of ASRU data via ESPNet. The PPL of the test data serves as the ground-truth baseline.
@@ V-A Comparison of model-centric methods
We further evaluated self-supervised learning (SSL) speech representations by training a Conformer-based ASR system. Results indicate that SSL speech representations significantly improve CS-ASR performance, which aligns with current research [32]. Specifically, multilingual features extracted from XLSR outperformed monolingual features extracted from WavLM. This suggests that multilingual pre-training provides more robust cross-lingual information that is beneficial for modeling code-switched speech.
@@
[…]
@@
Whisper-medium [43] - 12.1 87.1 62.6
+ $ft$ full model 762.3 M 6.8 11.5 16.5
Whisper-large-v3 [43] - 10.3 73.5 53.3
+ $ft$ full model 1541.6 M 6.6 10.3 14.9
@@ V-B Impact of Pre-training
When pre-trained models like Whisper are fine-tuned exclusively on code-switching data, the system degrades their performance on existing languages. This is a well-documented phenomenon in the literature [21]. It is often attributed to catastrophic forgetting or a distributional shift as the model adapts to achieve high performance on an unseen domain.
To explore solutions to the distributional shift, we also evaluated the impact of pre-training strategies on CS-ASR performance and presented the results in Table IV. Training a Conformer-L model directly on the mix of monolingual data and code-switching data obtains higher performance than sequentially pre-training on monolingual data and fine-tuning on code-switching data. This suggests that CS-ASR systems benefit from both cross-lingual and monolingual knowledge during training. While incorporating code-switching data in the training data moderately degrades monolingual ASR performance, it results in higher overall performance than both the pre-training and the fine-tuning strategies. The pre-trained Whisper-small model demonstrates an inherent ability to handle code-switching due to its large-scale multilingual pre-training. Fine-tuning the Whisper-small model on the ASRU leads to higher performance on both in-domain ASRU Mandarin and code-switching test data, while significantly degrading the performance on the LibriSpeech test-clean set.
@@
TABLE V: Comparison of fine-tuning the Whisper-small model on real, synthesized speech, and data augmented with synthesized speech. Here, the target texts and prompt samples are sourced from the original ASRU dataset. Real and synthesized data are denoted as “Real” and “Syn.”, respectively. “Speed Pert.” is speed perturbation
@@ V-C TTS with original text and speaker variety
Results indicate that augmenting training with synthesized speech and original text improves performance on SEAME but degrades performance on ASRU. When augmented with synthesized data three times the size of the original training set, the fine-tuned model achieves the best WER of 12.7% and 17.7% on the two SEAME test sets, respectively. However, this data augmentation strategy leads to performance degradation for ASRU. This discrepancy can be attributed to differences in textual complexity between the SEAME and ASRU datasets. We also compare this augmentation strategy with the traditional speed perturbation method. The results show that augmenting data with synthesized speech and original text yields higher performance than speed perturbation under the same data scale. Moreover, combining both augmentation methods leads to further performance gains on the SEAME dataset.
@@
V-D Impact of accents and language confusion on ASR
Since the $\text{dev}_{\texttt{man}}$ set has a CMI closer to the SEAME training data, the performance gap between $\text{dev}_{\texttt{man}}$ and $\text{dev}_{\texttt{sge}}$ presented in Table III suggests that the model performs worse on data with a higher degree of code mixing. To further examine how accent and language confusion in code-switching data affect ASR performance, we synthesized speech using reference prompts and target texts from multiple datasets, respectively. The resulting synthetic data were then used to fine-tune the Whisper-small model, and performance was compared across conditions as shown in Table VI.
To simulate accent mismatch, we paired text samples from the target dataset with prompt speech samples from a different dataset while keeping the TTS model consistent with the target dataset. For example, using a TTS model fine-tuned on SEAME data to synthesize speech for SEAME text with LibriSpeech or ASRU reference results in an accent mismatch. Conversely, the textual mismatch, primarily associated with syntactic and lexical characteristics, was simulated by pairing out-of-domain text with an in-domain prompt and TTS model.
@@ V-E Evaluation of the proposed SECT prompt
TABLE VII: Comparison of code-switching text synthesis with different LLMs and prompts, performance is evaluated with code-switching samples synthesized via 1500 ASRU Mandarin text samples, where the model is fine-tuned on 1500 synthesized speech-text pairs and tested on the ASRU test set. The CMI of the ASRU training set is 17.0, the PPL of the ASRU test set is 310.8
Prompting method ASR GPT-4o Human CMI PPL
@@
Baseline prompt 22.9 6.8 3.9 32.2 1794.5
EZSwitch prompt [20] 22.2 6.0 5.0 24.6 934.5
SECT prompt (ours) 19.2 7.0 6.7 17.3 540.2
w/o few-shot learning 19.6 7.0 5.7 15.2 564.5
@@
We evaluated the effectiveness of the proposed SECT-guided text generation method on 1,500 synthesized text samples in terms of WER for ASR performance, GPT-4o scores, and CMI. ASR performance was evaluated by fine-tuning a Whisper-small model on 1500 synthesized speech-text pairs before test, where TTS is done through the CosyVoice 2 model fine-tuned on the ASRU data. GPT-4o scores were obtained via an LLM-as-a-judge protocol, used together with CMI and human evaluation to assess the naturalness, grammatical correctness, and semantic preservation of the generated text samples. PPL were computed by a Transformer LM trained on the ASRU training set, which reflects the average uncertainty in next-token prediction.
The comparison between SECT-guided prompts and other prompting strategies is presented in Table VII. and the prompting strategies and LLM-as-a-judge evaluation are introduced in Table VIII. The results demonstrate that the SECT-guided prompting strategy achieves the best performance in terms of WER and GPT-4o scores, and yields CMI and PPL closest to that of the real ASRU code-switching text among all evaluated text generation strategies.
The PPL presented in Table VII indicates that our approach produces text that more closely aligns with the linguistic distribution of target code-switched speech. Given that the PPL values for both the test set and the generated text are significantly higher than that of the training data (4.3), we further verified this trend using a 3-gram word-level LM. The 3-gram LM achieves a PPL of 16.2 on the training data compared to 324.4 on the test data. This significant discrepancy is consistent with observations from the Transformer-based LM. These suggest that code-switching data inherently exhibits higher linguistic complexity and language confusion than monolingual data. The significantly high PPL of the generated text can likely be attributed to the presence of out-of-vocabulary (OOV) words and characters.
We also compared the LLM-as-a-judge with human evaluation in Table VII. Results show that the LLM-as-a-judge tends to overestimate the quality of text generated by LLM. In addition, human evaluation indicates that LLM-generated text does not yet fully replicate the linguistic distribution of spontaneous bilingual speech. Despite this gap, the proposed SECT method exhibits the highest score in the human evaluation, being consistent with LLM-as-a-judge. This demonstrates the effectiveness of the proposed SECT prompting strategy in code-switching text generation.
Moreover, we conducted ablation studies by comparing three prompting strategies: the baseline prompt, the SECT prompt without in-context learning, and the SECT prompt with in-context learning. Results confirm the effectiveness of the SECT-based prompting strategy, since it consistently outperforms the baseline regardless of whether in-context examples are included. Although incorporating in-context learning examples does not improve the GPT-4o scores, it leads to lower WER as well as CMI and PPL values that more closely approximate the real ASRU data. This implies that while ECT provides beneficial structural constraints for generating linguistically valid code-switching text, it may also impose rigid switching patterns. Incorporating in-context examples helps mitigate such rigidity, leading to more natural code-switching behavior.
@@ V-F TTS with text generated via SECT-prompted LLM
TABLE IX: Comparison between developing ASR models on ASRU data with different data augmentation methods. Text synthesis is performed via LLM guided by the proposed SECT prompts, and TTS is performed using target texts sourced from “Text source” and reference prompts sourced from the ASRU training set. Real and synthesized data are denoted as “Real” and “Syn.”, respectively
@@
+ ASRU Man. - 193+100 8.9
+ ASRU Man. - 193+500 9.2
+ Syn. w/ Real text ASRU 193+180 9.3
+ Syn. text w/ DeepSeek ASRU Mandarin 193+100 8.5
+ Syn. text w/ DeepSeek ASRU Mandarin 193+247 8.3
+ Syn. text w/ DeepSeek ASRU Mandarin 193+509 8.4
@@
+ $ft$ Syn. text w/ Qwen ASRU 352 18.4
+ $ft$ Syn. text w/ DeepSeek ASRU 509 13.8
Conformer-L ASRU - 193 11.8
+ Syn. text w/ DeepSeek ASRU Mandarin 193+247 10.2
Whisper-medium - - 12.1
+ $ft$ ASRU - 193 6.8
  + Syn. text w/ DeepSeek ASRU Mandarin 193+247 6.6
Whisper-large-v3 - - 10.3
+ $ft$ ASRU - 193 6.6
  + Syn. text w/ DeepSeek ASRU Mandarin 193+247 6.8
@@
The results in Table IX also demonstrate that the proposed SECT-based text generation method effectively augments the training of the Conformer-L model, yielding a 13.6% relative improvement. Regarding the Whisper models, the proposed method provides consistent performance gains for the small and medium sizes. However, a performance degradation is observed when fine-tuning Whisper-large-v3. This implies that for models with significantly high parameter counts, the specific distribution of the augmented data may induce overfitting. This consequently compromises the generalization capability of the model.
We also investigated the impact of varying the amount of synthesized speech–text pairs on CS-ASR performance. Results show that increasing the amount of pure Mandarin data during Whisper fine-tuning leads to performance degradation. This finding is consistent with the results presented in Table V, where data augmentation with limited code-switching patterns results in lower performance on ASRU data.
In contrast, incorporating synthesized code-switching speech–text pairs into fine-tuning yields moderate improvements. With a fixed baseline of 193 hours of real speech, we evaluated three data augmentation scales to investigate the relationship between real and synthetic data. In the low augmentation setting using 100 hours of synthetic speech, the model benefits moderately from the introduction of synthetic code-switching patterns. Increasing the synthetic data to 247 hours, a balanced augmentation, exhibits the highest performance. This is attributed to doubling the exposure to varied syntactic code-switching patterns while maintaining a strong anchor in real acoustic distributions. Model performance degrades when the volume of synthetic data is further increased to 509 hours, a state of synthetic over-representation. This suggests that the model starts to overfit to the specific acoustic artifacts and prosodic regularities of the TTS model. The resulting overfitting leads to a distribution shift, consequently compromising generalization on real-world human speech. Furthermore, this trend implies that the synthesized text data still falls short of real-world text in terms of naturalness and linguistic richness.
@@
TABLE X: A case study for error analysis of CS-ASR on SEAME dataset, with Conformer model, fine-tuned Whisper-small model, zero-shot Whisper-large-v3 model, and fine-tuned Whisper-large-v3 model
Method Output
Conformer not really ah sometime 我在家有自己帮兵
relax ah mm 你会 meet friends 吗对不对
还有 opportunity 就呃那个呃
Whisper-small ftft not really lah sometimes 我在家有自己帮人做
relax ah en 你会 meet friends 吗对不对
还有 opportunity 就是呃去那个呃
Whisper-large-v3 not really sometimes i will just pump
relax you will meet friends right
and the opportunity to go to
Whisper-large-v3 ftft not really lah sometimes 我在家有自己帮忙
relax ah en 你 会 meet friends mah 对不对
还有 opportunity 就是呃去那个呃
Groundtruth not really lah sometime 我在家里有自己 pumping
relax ah hum 因为 meet friends 嘛对不对
还有 opportunity 就是嗯去那个嗯
V-G Case study for error analysis
Existing research has explored how interjections, such as lah, ah, and lor in SEAME, impact the CS-ASR performance [28]. In our study, the CS-ASR performance can reach 6.6% MER for a fine-tuned Whisper-larget-v3 model, while the performance on two SEAME test sets remains high, being 10.3% and 14.9%, respectively. As models have shown promising performance on the ASRU data, we thus conducted a case study with representative examples as presented in Table X.
One common error involves phonetic confusion, where the English word, pumping, is incorrectly decoded as the Mandarin homophone. This is directly attributed to language confusion, as the decoder fails to leverage sufficient linguistic context to distinguish between the two languages at the switch point. Additionally, interjections remain a challenge for all models, which is consistent with findings in [28]. The examples also show that Whisper-large-v3 often exhibits a monolingual bias with a zero-shot setup, where it fails to switch and stays in a single language.
VI Discussion
VI-A Dynamic nature of code-switching
The dynamic nature of code-switching results in a combinatorial complexity. Unlike monolingual speech, code-switching does not adhere to a fixed vocabulary or a single grammar [39, 33]. Instead, it represents vast combinations of two or more languages. Treating this phenomenon as a third language ignores the nature of how speakers transition between languages. These make it inapplicable to synthesize code-switching text that aligns with an unknown target textual pattern.
Supported by established literature [61, 29, 58, 50, 53], our research suggests that utilizing monolingual data from the constituent languages significantly enhances CS-ASR performance. This trend remains consistent across different scenarios, such as the SEAME and ASRU datasets. These findings suggest that models benefit more from understanding the interaction between existing linguistic systems than from attempting to learn a new linguistic pattern from scratch.
VI-B Limitations of LLM-based TTS
In this study, we merely focused on fine-tuning the LLM module, which may result in limitations. In the CosyVoice 2 architecture, the LLM processes concatenated text and speech embeddings to model the conditional probability of output speech tokens. Restricting adaptation to the LLM module may lead to a cross-modal representation mismatch between the updated linguistic features and the fixed speech embeddings. This can thus limit the capability of the TTS model to achieve optimal acoustic naturalness and prosodic flow in the synthesized speech, particularly at code-switching boundaries.
VII Conclusion
In this paper, we investigated CS-ASR from both model- and data-centric perspectives, focusing on language confusion, accent bias, and data scarcity. Our model-centric analysis systematically reviewed language-aware algorithms, including language-specific processing and multi-task optimization. The evaluation indicates that these algorithms effectively mitigate language confusion. However, their performance is partially constrained by high accent bias, where language-discriminative information is inherently weakened. The data-centric analysis demonstrates that the optimal data augmentation strategy is highly dependent on the code-switching characteristics of the target domain. In high-frequency scenarios, textual patterns are naturally diverse. Therefore, the primary challenge is achieving acoustic robustness against rapid phonetic transitions. Standard augmentation methods are sufficient in these cases. Conversely, low-frequency scenarios exhibit an acoustic profile clearly biased toward the matrix language while textual variety remains inherently constrained. Consequently, the priority shifts toward increasing textual and structural diversity.